two private airplanes travel toward each other from cities that are 780 miles apart at speds 190 mph and 200mph. they left at the same time in how many hours will they meet? solution__________
Found 2 solutions by vleith, shahid:
Answer by vleith(2983)

Set up two equations for the 'position' of the two planes
P1 = 190t
P2 = 780 - 200t
Now find the time when the two Positions are equal
190t = 780 - 200t
390t = 780
t = 2
They will pass each other two hours into their flights

Answer by shahid(44)

as the aeroplanes are coming towards each other so their speeds will sum up to get the total speed so (200+190)=390 mph since distance is 780 miles now use formula distance= speed x time
780 = 390 x time
time = 780/390
time = 2 hours